SENSORS ACROSS THE SPECTRUM

\underline{CHRISTOPHER F. NEESE}, FRANK C. DE LUCIA, Department of Physics, The Ohio State University, 191 W. Woodruff Ave., Columbus, OH 43210 USA; IVAN R. MEDVEDEV, Department of Physics, Wright State University, 3640 Colonel Glenn Hwy, Dayton, OH 45435.

A resurgence of interest in spectroscopic sensors has been fueled by increases in performance made possible by technological advancements and applications in medicine, environmental monitoring, and national security. Often this research is technology driven, without enough consideration of the spectroscopic signatures available to be probed. We will compare several current spectroscopic sensors across the electromagnetic spectrum, with an eye towards the fundamental spectroscopic considerations important at each wavelength.